OpenStudy (anonymous):

"Given the position function of s of a point P moving on a coordinate line l, find the times at which the velocity is the given value k." s(t)=3t^(2/3) ; k=4 The answer in the book says it is 1/8. I am having trouble approaching this problem to get to the answer.

OpenStudy (anonymous):

Basically the velocity is the derivative: \[ v(t) = s'(t) \]And they want you to find the times (\(t\) values), where\[ v(t)=k \implies v(t)=4 \]

OpenStudy (anonymous):

So would \[s'(t)=2t ^{-2/3}\]

OpenStudy (anonymous):

First do \[ \frac 23-1 = \frac 23 - \frac 33=-\frac 13 \]

OpenStudy (anonymous):

Oh I see so it's \[2t ^{-1/3}\]

OpenStudy (anonymous):

You have to solve for \(t\):\[ 2t^{-1/3}=4 \]

OpenStudy (anonymous):

Ok so I see the derivative equal to 4 then solve for t. Now i stuck at \[t ^{-1/3}=2\] Hopefully I am solving this correctly.

OpenStudy (anonymous):

set*

OpenStudy (anonymous):

You aren't sure how to cancel out that exponent?

OpenStudy (anonymous):

I kinda forgot

OpenStudy (anonymous):

Need a little refresher

OpenStudy (anonymous):

Try taking both sides by the natural log of 2. \[ \log_2(t^{-1/3})=\log_2(2) \]

OpenStudy (anonymous):

wouldn't that just be end up as \[t ^{-1/3}=2\]

OpenStudy (anonymous):

It would give you \[ -\frac 13 \log_2(t)=1 \]

OpenStudy (anonymous):

ooh so you pull out the exponent to the front?

OpenStudy (anonymous):

Okay so when you simplify you get: \[ \log_2(t)=-3 \]

OpenStudy (anonymous):

The anti log of both sides: \[ t=2^{-3} \]

OpenStudy (anonymous):

So, as you can see, from the beginning all you really had to do with take both sides to the power of \(-3\)

OpenStudy (anonymous):

\[ (t^{-1/3})^{-3}=2^{-3} \]

OpenStudy (anonymous):

The reason I showed you the logarithm part is because it seems like your algebra is rusty and I wanted to teach logarithms.

OpenStudy (anonymous):

oh! omg i remember that now, that was so simple.can't believe i forgot. thanks a lot! :D

OpenStudy (anonymous):

yeah my algebra is verrrry rusty. i have to go back and review.